CALL FOR PAPERS

♢ Workshop on Argumentation and Nonmonotonic Reasoning (Arg-NMR) ♢

In conjunction with the Ninth International Conference on Logic Programming and Nonmonotonic Reasoning (LPNMR-07)

http://lia.deis.unibo.it/confs/ArgNMR/

Tempe, Arizona May 14, 2007

⇒ Aims and Scope ⇐

Research on Argumentation and Nonmonotonic Reasoning began in full force in the early eighties. The first attempts showed how argumentation results in a very natural way of conceptualizing Commonsense Reasoning, appropriately reflecting its defeasible nature. Further work in the KR&R community has shown that argumentation provides a useful perspective for relating different nonmonotonic formalisms. More recently, argumentation has been revealed as a powerful conceptual tool for exploring the theoretical foundations of reasoning and interaction in Autonomous Agents and Multiagent Systems.

This workshop will represent an opportunity for exchanging ideas on the fundamental theoretical basis and the design and implementation of argument-based systems including semantics, proof theory, applications to epistemic and practical reasoning, and the comparison of those systems with other types of nonmonotonic reasoning.

⇒ Topics ⇐

We solicit unpublished papers that present work on argumentation and nonmonotonic reasoning. We will privilege articles who emphasize connections between them. Relevant topics include, but are not limited to, the following:

□ Argumentation theories and logical foundations □ Argumentation and logic programming □ Formal models of argument □ Semantics of argumentation □ Operational semantics and execution models of

argumentation systems □ Argumentation and commonsense reasoning □ Argumentation for practical reasoning and

deliberation □ Argumentation tools & applications □ Argumentation for reasoning in multiagent systems □ Argumentation dialogues in multiagent systems □ Nonmonotonic reasoning in multiagent systems □ Argumentation for legal reasoning □ Argumentation and NMR in the semantic web □ Implementations of argumentation systems

⇒ Organization ⇐

□ Guillermo Simari, U Nac. del Sur, Bahia Blanca, Arg. □ Paolo Torroni, U Bologna, Italy

⇒ Submissions ⇐

We welcome and encourage the submission of high quality, original papers, which have not been accepted for publication nor are currently under review for another journal or conference. Papers should be written in English, formatted according to the Springer LNCS style (http://www.springer.com/comp/lncs/Authors.html), and they should not exceed sixteen (16) pages including title page, figures, references, etc.

⇒ Important Dates ⇐

□ Submission: 8 February 2007 □ Notification: 12 March 2007 □ Camera-ready: 29 March 2007 □ ArgNMR: 14 May 2007

⇒ Proceedings & Post-production ⇐

A printed volume with the proceedings will be available at the workshop. The proceedings of ArgNMR are also planned to form the basis for publishing a post-workshop volume, and/or a special issue of an international journal, subject to appropriate quality.

⇒ Programme Committee ⇐

□ Leila Amgoud, IRIT-CNRS Toulouse, France □ Grigoris Antoniou, FORTH-ICS, Greece □ Pietro Baroni, U Brescia, Italy □ Trevor Bench-Capon, U Liverpool, UK □ Carlos Chesñevar, U Nac. del Sur, Bahia Blanca, Arg. □ Jürgen Dix, TU Clausthal, Germany □ Phan Minh Dung, Asian Inst. of Technology, Thailand □ Lluis Godo, IIIA-CSIC, Spain □ Anthony Hunter, U College London, UK □ Antonis Kakas, U Cyprus □ Gabriele Kern-Isberner, U Dortmund, Germany □ Nicolas Maudet, U Paris-Dauphine, France □ Peter McBurney, U Liverpool, UK □ Donald Nute, U. Georgia, Athens, GE, US □ Henry Prakken, U Groeningen, U Utrecht, NL □ Iyad Rahwan, British U Dubai, UAE, U Edinburgh, UK □ Cao Tran Son, New Mexico State Univ., NM, US □ Francesca Toni, Imperial College London, UK